Two people, A and B, are sitting on a bench 0.70m apart. Person A has a mass of 60kg and person b experts a gravitational force of 8.0 x 10^-7 N on person a. Calculate the mass of person b.
1
Expert's answer
2019-04-30T09:58:03-0400

The gravitational force is


"F=\frac{Gm_am_b}{r^2}"

"m_b=\frac{Fr^2}{Gm_a}"

"m_b=\frac{(8\cdot 10^{-7})(0.7)^2}{(6.67\cdot 10^{-11})(60)}=98\ kg"